# RLAIF-V: Aligning MLLMs through Open-Source AI Feedback for Super GPT-4V Trustworthiness

### v1

@@ Appendix C Implementation Details
Table 9: Performance of different combine strategies. Acc.: Accuracy, MHuman.: MHumanEval.
Method Pairwise Object MHuman. AMBER
Accuracy HalBench
@@
AVG 69.0 15.3 8.4 50.0 81.9 86.7
REJ-P 58.0 27.1 13.9 53.4 78.1 84.9
REJ-C 70.0 13.3 7.5 41.8 79.9 85.9
@@ C.1 Different Combine Strategies
Besides scoring each response with the number of rejected claims (REJ-N), we also implement two more different ways to conduct this combination process: (1) average pooling (AVG) computes the mean of claim agreement scores $S_{i}=\sum_{j=1}^{m}{s_{c_{j}}[0]}$; (2) percentage of rejection (REJ-P) also counts the number $n_{rej}$ of claims that have $p_{no}>p_{yes}$ and $S_{i}=\frac{n_{rej}}{m}$. Comparison results of different combination methods are shown in Table 9. We observe that REJ-C obtains the best pairwise accuracy and achieves the most significant hallucination reduction on Object HalBench and MHumanEval.
@@ C.2 Response split and question generation
C.3 Self-rewarding Feedback Collection
@@
Appendix D Qualitative Results
@@
Appendix E Potential Impact and Limitations
@@
Table 11: Prompts for self-rewarding feedback collection.
@@
<user-question> {question} </user-question> <response> {answer} </response>
@@
<user-question> {question} </user-question> <response> {answer} </response>

### v2

@@ Appendix C Implementation Details
Method Agreement. Object MHuman. AMBER
HalBench
@@
REJ-P 83.3 27.1 13.9 53.4 78.1 84.9
REJ-C 96.7 13.3 7.5 41.8 79.9 85.9
Table 9: Performance of different combine strategies. Agreement.: Human agreement of the constructed pairs, MHuman.: MHumanEval.
@@ C.1 Different Combine Strategies
Besides scoring each response with the number of rejected claims (REJ-N), we also try to use percentage of rejection (REJ-P), which counts the number $n_{rej}$ of claims that have $p_{no}>p_{yes}$ and $S_{i}=\frac{n_{rej}}{m}$. Comparison results of different combination methods are shown in Table 9. We observe that REJ-C obtains better pairwise accuracy and achieves promising hallucination reduction on Object HalBench and MHumanEval.
@@ C.2 Response split and question generation
C.3 No divide-and-conquer Feedback Collection
@@
Appendix D Analysis on RLHF-V Dataset
D.1 Response Generation Model
RLHF-V [66] relies on human annotators identify and correct hallucinations, whereas RLAIF-V obtains feedback from open-source models without requiring human labor. The high cost of RLHF-V makes it challenging to provide correctional feedback for each model. We investigated the open-source dataset [1] of RLHF-V and found there are no responses generated by LLaVA 1.5 7B. We list the detailed proportions of responses generated by different MLLMs in the RLHF-V dataset in Table 10.
Model Proportion
Muffin [65] 38.7%
LLaVA 1.0 [34] 28.8%
Zephyr_MM 14.6%
InstructBLIP [10] 13.1%
Qwen-VL-Chat [4]  4.9%
Table 10: Proportions of responses generated by different MLLMs in the RLHF-V Dataset.
D.2 Hallucination Distribution
Upon reviewing the detailed evaluation of different MLLMs on MMHal Bench [56], as shown in Table 11, we found significant variation in the fine-grained hallucination score across models. Specifically, the correlations between Muffin [65] and LLaVA 1.5 7B, and between LLaVA 1.0 and LLaVA 1.5 7B, are even negative. Since the RLHF-V dataset primarily includes data from these two models, its effectiveness is significantly reduced due to the limited shared hallucination distribution.
Model Correlation Attribute Adversarial Comparison Counting Relation Environment Holistic Other
LLaVA 1.0 [34] -0.08 0.67 0.00 1.75 1.58 1.50 1.25 1.50 0.67
Muffin [65] -0.09 1.92 3.00 1.25 1.67 1.25 2.33 1.92 2.08
OmniLMM [46] 0.81 4.92 3.33 3.00 2.42 3.42 3.42 1.75 2.83
LLaVA 1.5 [33] 1.00 3.83 2.08 2.75 1.75 2.17 2.67 2.00 1.67
Table 11: Fine-grained hallucination scores of different MLLMs on MMHal Bench, and their correlation with LLaVA 1.5.
Appendix E Qualitative Results
@@
Appendix F Potential Impact and Limitations
@@
⟨user-question⟩ {question} ⟨/user-question⟩⟨response⟩ {answer} ⟨/response⟩
@@
⟨user-question⟩ {question} ⟨/user-question⟩⟨response⟩ {answer} ⟨/response⟩
@@
Table 13: Prompts for no divide-and-conquer feedback collection.